OpenStudy (anonymous):

Carla has 1 gallon of paint. She is going to pour it into a paint tray that measures 10 inches wide, 10 inches long, and 7 cm deep.

OpenStudy (ybarrap):

ya, then wut?

OpenStudy (anonymous):

A student runs 100 meters in 11 seconds. What is the speed of the student?

OpenStudy (ybarrap):

Speed is distance traveled divided by time taken to travel that distance. so we have 100/11. Make sense?

OpenStudy (anonymous):

okay

OpenStudy (ybarrap):

okay, make sense ? Or just okay? :)

OpenStudy (anonymous):

Carla has 1 gallon of paint. She is going to pour it into a paint tray that measures 10 inches wide, 10 inches long, and 7 cm deep. (1 gallon = 231 in3, 1 inch = 2.54 cm) Thanks

OpenStudy (ybarrap):

So do you want to know if that gallon of paint will fill the tray without spilling over then?

OpenStudy (anonymous):

Yes

OpenStudy (ybarrap):

Ok. let's figure out the volume of the tray: Length x Width x Height: 10 * 10 * 7 = ? That's the volume of the tray in cubic inches.

OpenStudy (anonymous):

140

OpenStudy (ybarrap):

But we need to convert that 7 cm to inches right. So 7 cm * 1 inch/2.54 cm = 7/2.54 = depth of tray in inches.

OpenStudy (ybarrap):

So the volume of the tray is really, 10 * 10 *7/2.54 = ? This will then be in cubic inches.

OpenStudy (anonymous):

55.11?

OpenStudy (anonymous):

Which of the following scenarios will happen? The paint will not fill the tray by 44.59 in3. The paint will not fill the tray by 469 cm3. The paint will fill the tray exactly. The paint will overfill the tray by 44.59 in3.

OpenStudy (ybarrap):

Compare that to the volume of a gallon, which is 231 cubic inches. Is 10*10*7/2.54 = 276 greater or less than a gallon of paint?

OpenStudy (anonymous):

greater

OpenStudy (ybarrap):

By how much?

OpenStudy (anonymous):

44.59

OpenStudy (ybarrap):

So. Which choice makes most sense to you?

OpenStudy (anonymous):

D

OpenStudy (ybarrap):

I agree with you.

OpenStudy (anonymous):

Thank you can you help me with 1 more problem please

OpenStudy (ybarrap):

yep. ok

OpenStudy (anonymous):

Kenya is reserving a room in a hotel in France where they use euros (€). The room charge is €75. Suppose the conversion rate is €1 = $1.298. What is the cost of the room in dollars? $57.78 $77.56 $97.35 $129.80

OpenStudy (ybarrap):

For any type of conversion from one unit to the other use this trick: \( \cfrac{$1.29}{€1}\times€75 = \cfrac{$1.29}{\cancel{€}1}\times\cancel{€}75= \)? Notice how the units I'm converting FROM canceled and leaves me with the units I'm converting TO.

OpenStudy (ybarrap):

You are left with just 1.29 x 75 = ? In dollars.

OpenStudy (anonymous):

I got c

OpenStudy (ybarrap):

I agree with you again.

OpenStudy (anonymous):

Thanks for everything (;

OpenStudy (ybarrap):

yw